# Time Non-locality in Dark Matter and LSS

### v1

@@ Abstract
We explore the intriguing phenomenon of time non-locality in the evolution of dark matter and Large Scale Structure (LSS). Recently in [1], it was shown that time non-locality emerges in bias tracer fluctuations, which are scalars in real space, at fifth order in the perturbation expansion in dark matter overdensity. We demonstrate that by breaking the symmetry down to , which is the case whenever line-of-sight effects become important, such as for flux fluctuations in the Lyman forest, the temporal non-locality appears at the third order in expansion. Additionally, within the framework of EFTofLSS, we demonstrate that time non-locality manifests in the effective stress tensor of dark matter, which is a second rank tensor under transformations, again at the third order in dark matter overdensity. Furthermore, we highlight the effectiveness of the standard basis [2] in handling time non-local operators.
@@ 1 Introduction
The Standard Model of cosmology includes an inflationary phase of accelerated expansion, leading to homogeneity and isotropy on large scales, while at the same time giving rise to small inhomogeneities from quantum fluctuations. These inhomogeneities are reflected in the anisotropy of the cosmic microwave background (CMB) [3, 4, 5] , and also serve as initial density perturbations that grow during matter domination era and give rise to large scale structures (LSS). The evolution of LSS perturbations, therefore, contains information about the initial spectrum of fluctuations but is also sensitive to the relative ratios and properties of energy components that drive the subsequent expansion, including Dark Energy (DE) and Dark Matter (DM). Understanding the distribution of LSS is fundamental to our understanding of very early Universe physics, as well to discern the ultimate fate of the Universe.
@@
Due to the three-dimensional nature, the number of observable modes in LSS is larger compared to the CMB and therefore is, in principle, statistically more constraining on the parameters of interest. However, since gravitational evolution drives the perturbations into the nonlinear regime on small scales, accurate theoretical modeling down to these scales remains a challenge. The equations governing the dynamics of dark matter are the Boltzmann equations in an expanding background. One approach has been to use computationally intensive -body simulations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] of the dark matter phase space, under the assumption that DM is the major clustering component, subsequently driving the evolution of the baryon clustering. These studies have been extended beyond CDM, see for e.g. [19, 20, 21]. In a (hypothetical) Universe with DM as the only clustering component, these simulations are accurate down to the scales much smaller than the highly nonlinear, virialized halo structures.
@@
The second approach to treating gravitational clustering beyond linear order is to use higher-order perturbation theory. In the Standard Perturbation Theory [22, 23, 24, 25, 26, 27], the relevant equations are solved order by order using a series ansatz. However, even for quasilinear wavenumbers where the perturbation expansion should be technically valid, this approach is unable to reproduce the matter power spectrum computed through simulations. Therefore, principles of Effective Field Theories (EFT) have been applied to the context of the LSS perturbations to model the quasi-linear scales [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In this approach, termed the EFT of LSS (EFTofLSS), we integrate out the short scale modes in SPT. The process of integrating out short modes generates an effective stress tensor in the equations of motion. The stress tensor incorporates all the short-scale physics which is needed to describe the distribution of dark matter at quasi-linear scales. Various correlation functions of the density fluctuations can then be calculated within the framework. Predictions of EFTofLSS have been benchmarked against full simulations to a high degree of accuracy on quasilinear scales, and these techniques have now been applied to various datasets to extract state-of-the-art constraints on various cosmological parameters [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
In the perturbative approach, baryonic matter such as galaxies are assumed to be tracers of the underlying dark matter field. The overdensity for tracers is given as a “bias expansion" [53], in terms of the dark matter fields. In most cases, the expansion is assumed to be local in time - i.e. the overdensity of tracers at some time is written in terms of dark matter operators at the same time. However, it has been argued that since galaxies are not formed instantaneously — the formation time is of order of the evolution timescales of the dark matter field itself — the bias expansion should be non-local in time [54]. Recent investigations [1] have shown that, indeed, time non-locality is an essential feature of the bias expansion at the fifth order. Technically, it was shown that when writing a bias expansion consisting of zero derivative invariant operators of dark matter fields in real space, the local and the non-local basis differ at the fifth order in the perturbative expansion. This indicates that the physics of biased tracers cannot be described accurately by local expansion beyond the fourth order in perturbation theory.
@@
While the bias expansion for the clustering of tracers in real space is written in terms of invariant operators, there are observationally relevant cases in cosmology when additional invariant terms must also be included in the expansion. Such cases typically arise when the line-of-sight (LoS) direction becomes special, therefore breaking the symmetry. This is exactly what happens while considering the clustering of tracers in redshift space [55] — the main observable of spectroscopic surveys such as DESI [56]. Another case where such operators are essential is in the bias expansion of the flux fluctuations in the Lyman- forest [57]. It is, therefore, interesting to investigate the issue of time non-locality for such invariant operators. This is the first such study, and we find that time non-locality appears at the third order for these operators, as opposed to the the fifth order for the invariant operators. This implies that time non-locality may be detected in the 1-loop power spectrum for observables for which selection operators need to be included. In contrast, for the bias expansion, time non-locality can only be detected at the two-loop power spectrum of the observables.
@@
Another question that arises is whether the issue of time non-locality plays any role in the evolution of the dark matter field itself within the context of EFTofLSS. For the EFT approach to work, a separation of scales is needed between the scales at which we want predictions (i.e. the quasilinear scales) and the scales at which new (nonlinear) physics kicks in. The wavenumber at which the density perturbations become of (1) is known as the nonlinear wavenumber (). This, therefore, defines a natural scale in the space domain and makes it possible to define an effective field theory (EFTofLSS), by integrating out the modes above the non-linear scale. The presence of a scale separation in space ensures that integrating out the short modes gives a space local effective theory. The local EFT has a derivative expansion that is regulated by the non-linear scale (). However, the typical timescales associated with the evolution of both the large scales and the small scales, are similar . Therefore, interestingly, there is no scale separation in the time domain. Therefore, integrating out the short modes generically yields a time non-local EFT. This time non-locality manifests itself in the governing equations of LSS through the effective stress tensor in the equations of motion for dark matter [32, 58, 33, 34, 36]. The non-locality in stress tensor is because there is no hierarchy in evolution timescales between the long and the short modes. We have investigated the structure of the time non-local stress tensor beyond linear order. We have found that it differs from the time local stress tensor at third order. This is due to the tensorial nature of the stress tensor —- for scalars such as the difference between the local-in-time and non-local-in-time expressions is at the fourth order.
@@
The paper is organised as follows. In Sec. 2, we review the basic equations of SPT and biased tracers relevant to our discussion. We describe an alternate system (-system) of writing operators in the bias expansion and also discuss the structure of selection operators. We use this section to also set up the conventions and notations that will be used throughout the paper. In Sec. 3, we discuss the EFTofLSS and the effective stress tensor. Then, we discuss how time non-locality appears in dark matter and lay out the details of getting the non-local stress tensor up to third order. Then, in Sec. 4, we discuss time non-locality in the bias expansion of invariant operators. We also discuss the invertible mapping between the non-local and -basis. In Sec. 5, we discuss bias expansion with the invariant (selection) operators and the time non-local basis upto fourth order. Finally, in Sec. 6, we conclude by stating our results and giving some future directions. In appendices A and B we discuss the vector operator structure for dark matter and renormalisation of the flux field in Lyman respectively.
@@ 2.1 Dark-Matter: Standard Perturbation Theory (SPT)
Dark matter at large scales has been shown to behave as an effective fluid following the Vlasov-Poisson equation which is the dynamical equation describing collisionless matter coupled with gravity [28, 29]. In the single stream approximation i.e. for a vanishing velocity dispersion, the Boltzmann hierarchy can be truncated after the first moment and we end up with the continuity and Euler equation for the dark matter overdensity and velocity field. The continuity and Euler equation for collisionless fluid is given as
@@
2.2 Bias-Expansion
Baryonic matter, such as galaxies, traces the underlying dark matter field as a result of which the number density of these "tracers" is correlated with the dark matter overdensity. Due to the absence of an equation of motion for tracers, there is no first principle way to get the tracer overdensity. However, we know that the tracer overdensity can be affected by dark matter fields only locally and therefore, can only depend on the locally measurable quantities constructed out of the dark matter fields. Due to rotational invariance and the equivalence principle, the locally measurable quantities are the tidal field () and velocity gradient (). Tracer overdensity is a scalar quantity, therefore it can depend only on scalars constructed out of the tidal field and velocity gradient. In general, the relation between galaxy number overdensity and dark matter overdensity can be expressed with the help of a function such that the galaxy overdensity () is given as [54, 59],
where is the number density of galaxies and the bar represents the mean values. The function is such that it vanishes when all the arguments vanish. It contains higher derivative as well as stochastic terms, which we have suppressed. We can Taylor expand Eq.(2.5) about zero and get the following expansion,
where ’s are time-dependent coefficients also called "bias" coefficients which can be measured from simulation or galaxy data by galaxy-galaxy correlation or galaxy-dark matter cross correlations. We have just written the leading order terms in Eq.(2.6) and not the terms containing higher derivatives which are also present111In this work we will not be considering higher derivative and stochastic terms in the bias expansion for tracers.. We can write Eq.(2.6) in a more familiar and compact form as,
@@ 2.3 Selection operators
Physically, these selection operators are crucial while comparing the theoretical predictions with observations. This is due to the fact that we observe galaxies through the light that reaches us. The probability that we will observe a photon emitted from a galaxy in a particular direction depends on the line-of-sight of observation. These effects can be incorporated in the bias expansion by including new line of sight dependent operators [55] which are precisely the selection operators given in Eq.(2.14). In the context of galaxies, a subset of these selection operators appear when we move to galaxy clustering in redshift space.
However, in other contexts such as when writing an EFT for the flux field from Lyman- forest, we need to include all possible selection operators [57] in the bias expansion of the flux fluctuations. This happens due to the non-linear mapping between the flux field () and the optical depth (), which is a biased tracer of the underlying dark matter field. The non-linear mapping is given by an exponential as . Owing to this exponential map, a "renormalised" EFT of flux overdensity field, , with the optical depth taken to be in redshift space, contains all possible selection operators. We have commented on this in a bit more detail in Appendix. B. This way of generating all possible selection operators is termed as the "top-down" approach to writing EFTs as opposed to the "bottom-up" approach that we have discussed so far.
@@
In this section, we describe an alternate system of writing operators constructed out of long wavelength dark matter fields, the system [2]. However, the effectiveness of system will be shown in capturing the time non-locality in biased tracers or dark matter as discussed later in Secs. 4 and 3 respectively. The usual method to compute the set of independent non-local operators for biased tracers is discussed in [60, 1]. This method is very tedious and contains a lot of redundancies. However, we point out that when working with the system [2], the calculation of determining the set of independent non-local operators is much more direct and efficient.
For biased tracers of the dark matter field, such as halos and galaxies, [2] proposed a system for writing the relevant operators under the assumption that galaxy overdensity field follows the same continuity equation as the dark matter overdensity. We will use this system to write down the EFT operators appearing in the effective stress tensor relevant for the dark matter field itself and also to write the operators appearing in the bias expansion for tracers. The reason to do this is that they have a nice connection with time non-locality which will be discussed in later sections. In this system we define
We also define [2],

### v2

@@ Abstract
We explore the intriguing phenomenon of time non-locality in the evolution of dark matter and Large Scale Structure (LSS). Recently in [Donath:2023sav], it was shown that time non-locality emerges in bias tracer fluctuations, which are scalars in real space, at fifth order in the perturbation expansion in dark matter overdensity. We demonstrate that by breaking the symmetry down to , which is the case whenever line-of-sight effects become important, such as for flux fluctuations in the Lyman forest, the temporal non-locality appears at the third order in expansion. Additionally, within the framework of EFTofLSS, we demonstrate that time non-locality manifests in the effective stress tensor of dark matter, which is a second rank tensor under transformations, again at the third order in dark matter overdensity. Furthermore, we highlight the effectiveness of the standard basis [Mirbabayi:2014zca] in handling time non-local operators.
@@ 1 Introduction
The Standard Model of cosmology includes an inflationary phase of accelerated expansion, leading to homogeneity and isotropy on large scales, while at the same time giving rise to small inhomogeneities from quantum fluctuations. These inhomogeneities are reflected in the anisotropy of the cosmic microwave background (CMB) [WMAP:2012nax, Planck:2018vyg, Planck:2018jri] , and also serve as initial density perturbations that grow during matter domination era and give rise to large scale structures (LSS). The evolution of LSS perturbations, therefore, contains information about the initial spectrum of fluctuations but is also sensitive to the relative ratios and properties of energy components that drive the subsequent expansion, including Dark Energy (DE) and Dark Matter (DM). Understanding the distribution of LSS is fundamental to our understanding of very early Universe physics, as well to discern the ultimate fate of the Universe.
@@
Due to the three-dimensional nature, the number of observable modes in LSS is larger compared to the CMB and therefore is, in principle, statistically more constraining on the parameters of interest. However, since gravitational evolution drives the perturbations into the nonlinear regime on small scales, accurate theoretical modeling down to these scales remains a challenge. The equations governing the dynamics of dark matter are the Boltzmann equations in an expanding background. One approach has been to use computationally intensive -body simulations [Angulo:2021kes, hockney2021computer, Springel:2006vs, Vogelsberger:2019ynw, Chisari:2011iq, Adamek:2013wja, Adamek:2016zes, Adamek:2017grt, Ali-Haimoud:2012fzp, Angulo:2016hjd, Bayer:2020tko, Chudaykin:2020hbf, Foreman:2015lca] of the dark matter phase space, under the assumption that DM is the major clustering component, subsequently driving the evolution of the baryon clustering. These studies have been extended beyond CDM, see for e.g. [Banerjee:2016zaa, Banerjee:2018bxy, Banerjee:2019bjp]. In a (hypothetical) Universe with DM as the only clustering component, these simulations are accurate down to the scales much smaller than the highly nonlinear, virialized halo structures.
@@
The second approach to treating gravitational clustering beyond linear order is to use higher-order perturbation theory. In the Standard Perturbation Theory [Bouchet:1992xz, Catelan:1994kt, Matsubara:1995kq, Bouchet:1994xp, Scoccimarro:2000sn, Bernardeau:2001qr], the relevant equations are solved order by order using a series ansatz. However, even for quasilinear wavenumbers where the perturbation expansion should be technically valid, this approach is unable to reproduce the matter power spectrum computed through simulations. Therefore, principles of Effective Field Theories (EFT) have been applied to the context of the LSS perturbations to model the quasi-linear scales [Baumann:2010tm, Carrasco:2012cv, Porto:2013qua, Senatore:2014via, Carrasco:2013mua, Angulo:2014tfa, Baldauf:2014qfa, Konstandin:2019bay, Bertolini:2016bmt, Abolhasani:2015mra]. In this approach, termed the EFT of LSS (EFTofLSS), we integrate out the short scale modes in SPT. The process of integrating out short modes generates an effective stress tensor in the equations of motion. The stress tensor incorporates all the short-scale physics which is needed to describe the distribution of dark matter at quasi-linear scales. Various correlation functions of the density fluctuations can then be calculated within the framework. Predictions of EFTofLSS have been benchmarked against full simulations to a high degree of accuracy on quasilinear scales, and these techniques have now been applied to various datasets to extract state-of-the-art constraints on various cosmological parameters [Cabass:2022wjy, Cabass:2022ymb, Philcox:2021kcw, Ivanov:2021kcd, DAmico:2019fhj, Braganca:2023pcp, DAmico:2022osl, DAmico:2022gki, Zhang:2021yna, DAmico:2020tty, DAmico:2020kxu, Nishimichi:2020tvu, Colas:2019ret, Ivanov:2023qzb, Schmidt:2020viy].
@@
In the perturbative approach, baryonic matter such as galaxies are assumed to be tracers of the underlying dark matter field. The overdensity for tracers is given as a “bias expansion" [Desjacques:2016bnm], in terms of the dark matter fields. In most cases, the expansion is assumed to be local in time - i.e. the overdensity of tracers at some time is written in terms of dark matter operators at the same time. However, it has been argued that since galaxies are not formed instantaneously — the formation time is of order of the evolution timescales of the dark matter field itself — the bias expansion should be non-local in time [Senatore:2014eva]. The "local in time" and "non-local in time" terminology used here has been used in various published works [Donath:2023sav, Ivanov:2023yla, Carroll:2013oxa, Angulo:2015eqa, DAmico:2022ukl]. Hence, we use the same terminology to be consistent with the literature. A more formal definition of time local and time non- local operators will be given in later sections. Recent investigations [Donath:2023sav] have shown that, indeed, time non-locality is an essential feature of the bias expansion at the fifth order. Technically, it was shown that when writing a bias expansion consisting of zero derivative invariant operators of dark matter fields in real space, the local and the non-local basis differ at the fifth order in the perturbative expansion. This indicates that the physics of biased tracers cannot be described accurately by local expansion beyond the fourth order in perturbation theory.
@@
While the bias expansion for the clustering of tracers in real space is written in terms of invariant operators, there are observationally relevant cases in cosmology when additional invariant terms must also be included in the expansion. Such cases typically arise when the line-of-sight (LoS) direction becomes special, therefore breaking the symmetry. This is exactly what happens while considering the clustering of tracers in redshift space [Desjacques:2018pfv] — the main observable of spectroscopic surveys such as DESI [DESI:2024mwx]. Another case where such operators are essential is in the bias expansion of the flux fluctuations in the Lyman- forest [Ivanov:2023yla]. It is, therefore, interesting to investigate the issue of time non-locality for such invariant operators.
Time non-locality in an inherent part of the EFT framework for the reasons discussed above. This has motivated a lot of studies in this direction, see for example [Abolhasani:2015mra, Senatore:2014eva, Ivanov:2023yla, Carroll:2013oxa, Angulo:2015eqa, DAmico:2022ukl, Mirbabayi:2014zca, Desjacques:2016bnm] and references therein. As stated before, most recent study in this direction investigated invariant operators and found that time non-locality manifests itself at fifth order for such operators [Donath:2023sav]. Ours is first study which investigates time non-locality in the context of invariant operators, and we find that time non-locality appears at the third order for these operators, as opposed to the fifth order for the invariant operators. This implies that time non-locality may be detected in the 1-loop power spectrum for observables for which selection operators need to be included. In contrast, for the bias expansion, time non-locality can only be detected at the two-loop power spectrum of the observables.
@@
Another question that arises is whether the issue of time non-locality plays any role in the evolution of the dark matter field itself within the context of EFTofLSS. Indeed it is the case with EFTofLSS and has been discussed in the literature [Abolhasani:2015mra, Mirbabayi:2014zca, Desjacques:2016bnm]. Let us briefly explain the reason behind time non-locality in EFTofLSS. For the EFT approach to work, a separation of scales is needed between the scales at which we want predictions (i.e. the quasilinear scales) and the scales at which new (nonlinear) physics kicks in. The wavenumber at which the density perturbations become of (1) is known as the nonlinear wavenumber (). This, therefore, defines a natural scale in the space domain and makes it possible to define an effective field theory (EFTofLSS), by integrating out the modes above the non-linear scale. The presence of a scale separation in space ensures that integrating out the short modes gives a space local effective theory. The local EFT has a derivative expansion that is regulated by the non-linear scale (). However, the typical timescales associated with the evolution of both the large scales and the small scales, are similar . Therefore, interestingly, there is no scale separation in the time domain. Therefore, integrating out the short modes generically yields a time non-local EFT. This time non-locality manifests itself in the governing equations of LSS through the effective stress tensor in the equations of motion for dark matter [Carrasco:2013mua, Carroll:2013oxa, Angulo:2014tfa, Baldauf:2014qfa, Bertolini:2016bmt]. The non-locality in stress tensor is because there is no hierarchy in evolution timescales between the long and the short modes. We have investigated the structure of the time non-local stress tensor beyond linear order. The time local expansions offer a simpler, and physically intuitive framework. Given its simplicity, and the fact that the two expansions match at the leading orders, the time local expansion is a useful tool for practical calculations. Our objective in this work is to point out up to what order in perturbation theory one can write time local counterterms in the effective stress tensor and still be robust in the analysis. We have found that the time non-local stress tensor differs from the time local stress tensor at third order. This is due to the tensorial nature of the stress tensor —- for scalars such as the difference between the local-in-time and non-local-in-time expressions is at the fourth order, see Appendix. B.
We have demonstrated that the relevance of time non-local effects at a given order depends on the specific observable under consideration. For instance, when examining observables that depend solely on invariant operators in the bias expansion, time non-local effects manifest at fifth order [Donath:2023sav]. Conversely, for observables that also depend on projections of dark matter fields (i.e. invariant structures), time non-local effects become significant at third order. Clarifying this distinction is also an objective of our analysis.
@@
The paper is organised as follows. In Sec. 2, we review the basic equations of SPT and biased tracers relevant to our discussion. We describe an alternate system (-system) of writing operators in the bias expansion and also discuss the structure of selection operators. We use this section to also set up the conventions and notations that will be used throughout the paper. In Sec. 3, we discuss the EFTofLSS and the effective stress tensor. Then, we discuss how time non-locality appears in dark matter and lay out the details of getting the non-local stress tensor up to third order. Then, in Sec. 4, we discuss time non-locality in the bias expansion of invariant operators. We also discuss the invertible mapping between the non-local and -basis. In Sec. 5, we discuss bias expansion with the invariant (selection) operators and the time non-local basis upto fourth order. Finally, in Sec. 6, we conclude by stating our results and giving some future directions. In appendices A and C we discuss the vector operator structure for dark matter and renormalisation of the flux field in Lyman respectively.
@@ 2.1 Dark-Matter: Standard Perturbation Theory (SPT)
Dark matter at large scales has been shown to behave as an effective fluid following the Vlasov-Poisson equation which is the dynamical equation describing collisionless matter coupled with gravity [Baumann:2010tm, Carrasco:2012cv]. In the single stream approximation i.e. for a vanishing velocity dispersion, the Boltzmann hierarchy can be truncated after the first moment and we end up with the continuity and Euler equation for the dark matter overdensity and velocity field. The continuity and Euler equation for collisionless fluid is given as
@@
2.2 Time local Bias Expansion
Baryonic matter, such as galaxies, traces the underlying dark matter field as a result of which the number density of these "tracers" is correlated with the dark matter overdensity. Due to the absence of an equation of motion for tracers, there is no first principle way to get the tracer overdensity. However, we know that the tracer overdensity can be affected by dark matter fields only locally and therefore, can only depend on the locally measurable quantities constructed out of the dark matter fields. Due to rotational invariance and the equivalence principle, the locally measurable quantities are the tidal field () and velocity gradient (). Tracer overdensity is a scalar quantity, therefore it can depend only on scalars constructed out of the tidal field and velocity gradient.
This allows us to write the tracer overdensity as follows [Senatore:2014eva, McDonald:2009dh],
where , being the number density of galaxies and ’s are time-dependent coefficients also called "bias" coefficients which can be measured from simulation or galaxy data by galaxy-galaxy correlation or galaxy-dark matter cross correlations. We have just written the leading order terms in Eq.(2.5) and not the terms containing higher derivatives which are also present111In this work we will not be considering higher derivative and stochastic terms in the bias expansion for tracers.. Note that Eq.(2.5) is not the most general form for tracer overdensity since it is local-in-time. The most general form for tracer overdensity is time non-local and will be discussed later in Sec. 4. We can write Eq.(2.5) in a more familiar and compact form as,
@@ 2.3 Selection operators
Physically, these selection operators are crucial while comparing the theoretical predictions with observations. This is due to the fact that we observe galaxies through the light that reaches us. The probability that we will observe a photon emitted from a galaxy in a particular direction depends on the line-of-sight of observation. These effects can be incorporated in the bias expansion by including new line of sight dependent operators [Desjacques:2018pfv] which are precisely the selection operators given in Eq.(2.13). In the context of galaxies, a subset of these selection operators appear when we move to galaxy clustering in redshift space.
However, in other contexts such as when writing an EFT for the flux field from Lyman- forest, we need to include all possible selection operators [Ivanov:2023yla] in the bias expansion of the flux fluctuations. This happens due to the non-linear mapping between the flux field () and the optical depth (), which is a biased tracer of the underlying dark matter field. The non-linear mapping is given by an exponential as . Owing to this exponential map, a "renormalised" EFT of flux overdensity field, , with the optical depth taken to be in redshift space, contains all possible selection operators. We have commented on this in a bit more detail in Appendix. C. This way of generating all possible selection operators is termed as the "top-down" approach to writing EFTs as opposed to the "bottom-up" approach that we have discussed so far.
@@
In this section, we describe an alternate system of writing operators constructed out of long wavelength dark matter fields, the system [Mirbabayi:2014zca]. However, the effectiveness of system will be shown in capturing the time non-locality in biased tracers or dark matter as discussed later in Secs. 4 and 3 respectively. The usual method to compute the set of independent non-local operators for biased tracers is discussed in [DAmico:2022ukl, Donath:2023sav]. This method is very tedious and contains a lot of redundancies. However, we point out that when working with the system [Mirbabayi:2014zca], the calculation of determining the set of independent non-local operators is much more direct and efficient.
For biased tracers of the dark matter field, such as halos and galaxies, [Mirbabayi:2014zca] proposed a system for writing the relevant operators which stems from the fact that convective derivative of a locally observable operator is also locally observable [Mirbabayi:2014zca, Desjacques:2016bnm]. We will use this system to write down the EFT operators appearing in the effective stress tensor relevant for the dark matter field itself and also to write the operators appearing in the bias expansion for tracers. The reason to do this is that they have a nice connection with time non-locality which will be discussed in later sections. In this system we define
We also define [Mirbabayi:2014zca],